# MCMC Marginalisation Bias and $\Lambda$CDM tensions

### v2

@@ Acknowledgments
We would like to thank Adrià Gómez-Valent for discussions and comments on the draft. We thank Gabriela Marques, Mike Hudson and Matteo Viel for related discussions on late Universe evolution in $S_{8}$. EÓC thanks Yonsei University and Asia Pacific Center for Theoretical Physics for hospitality. This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology). MMShJ thanks the support from ICTP associates office (under Senior Associate program) and ICTP HECAP section for hospitality. SP and MMShJ acknowledge SarAmadan grant No. ISEF/M/401332.

### v3

@@ Acknowledgments
We thank Adrià Gómez-Valent and Leandros Perivolaropoulos for discussions and comments on the draft. We thank Gabriela Marques, Mike Hudson and Matteo Viel for related discussions on late Universe evolution in $S_{8}$. EÓC thanks Yonsei University and Asia Pacific Center for Theoretical Physics for hospitality. This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology). MMShJ thanks the support from ICTP associates office (under Senior Associate program) and ICTP HECAP section for hospitality. SP and MMShJ acknowledge SarAmadan grant No. ISEF/M/401332.